# SIM2VR: Towards Automated Biomechanical Testing in VR

### v1

@@ Abstract.
Automated biomechanical testing has great potential for the development of VR applications, as initial insights into user behaviour can be gained in silico early in the design process. In particular, it allows prediction of user movements and ergonomic variables, such as fatigue, prior to conducting user studies. However, there is a fundamental disconnect between simulators hosting state-of-the-art biomechanical user models and simulators used to develop and run VR applications. Existing user simulators often struggle to capture the intricacies and nuances of real-world VR applications, reducing ecological validity of user predictions. In this paper, we introduce SIM2VR, a system that aligns user simulation with a given VR application by establishing a continuous closed loop between the two processes. This, for the first time, enables training simulated users directly in the same VR application that real users interact with. We demonstrate that SIM2VR can predict differences in user performance, ergonomics and strategies in a fast-paced, dynamic arcade game. In order to expand the scope of automated biomechanical testing beyond simple visuomotor tasks, advances in cognitive models and reward function design will be needed.
biomechanical simulation, interaction design, VR simulation alignment problem, automated testing, virtual reality, VR development, VR application, deep reinforcement learning
copyright: acmcopyrightjournalyear: 2018copyright: rightsretainedconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYbooktitle: Woodstock ’18: ACM Symposium on Neural Gaze Detection, June 03–05, 2018, Woodstock, NYdoi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06ccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Systems and tools for interaction designccs: Human-centered computing Virtual realityccs: Human-centered computing User models
Figure 1. Simulating interactive user behaviour in VR typically requires reimplementing the game dynamics, logic and graphics of the VR application in another simulation engine, which is time-consuming, redundant and error-prone (left).
SIM2VR (middle) enables VR designers for the first time to run state-of-the-art biomechanical user simulations directly in the same VR application real users are interacting with (right). By aligning simulated and real interaction with respect to VR input and output, as well as application dynamics, SIM2VR allows for more ecologically valid predictions of user performance, effort and strategies, supporting developers in the early stages of the design process.
@@ 1. Introduction
Recent advances in deep reinforcement learning (deep RL) and biomechanical models have enabled the creation of simulated users for modelling how humans may behave in interaction tasks. Using advanced biomechanical models that capture the intricacies of human body movement beyond linked-segment models, such simulated users allow prediction of time-continuous movements and effort-related quantities such as muscle energy expenditure (Umberger et al., 2003) and fatigue (Xia and Law, 2008; Rashedi and Nussbaum, 2015). In HCI, simulated users have been recently demonstrated for various tasks, including pointing (Klar et al., 2023; Cheema et al., 2020), typing (Hetzel et al., 2021; Jokinen et al., 2021), tracking (Fischer et al., 2021), steering and choice reaction (Ikkala et al., 2022). This modeling approach is promising, as it enables designers and developers to pose “what-if?” questions to assess how changes in design choices affect users’ performance and ergonomics.
@@
This bears particular promise for HCI domains where human data is scarce or costly to collect, such as in virtual reality (VR) interactions. In VR development, testing prototypes typically involves wearing a headset, standing up and moving around in space, which incurs significant overhead and renders iterative testing slow and cumbersome (Liu et al., 2023). Biomechanical user simulations show significant potential to remedy this as a means of automated testing. In particular, they allow VR designers to evaluate and compare prototypes without the need for expensive user studies, which can be lengthy and stressful, especially for vulnerable participants. By using an appropriate range of different user models, a simulation-based approach can increase diversity during the testing process, thereby promoting ability-based design (Murray-Smith et al., 2022).
@@
To realise the potential of simulated users, we need to address the challenge of integrating model-based simulations into VR development workflows. State-of-the-art biomechanical models are not readily available in commonly used VR development environments like Unity [^1] https://unity.com/, or they cannot be combined with deep RL approaches to generate forward simulations in those environments. For instance, while there exists a MuJoCo Unity plugin [^2] https://github.com/google-deepmind/mujoco/tree/main/unity that provides access to biomechanical models implemented in MuJoCo (Todorov et al., 2012), an efficient physics engine suitable for biomechanical simulation, the models can only be controlled manually using the Unity Editor.
@@
As a result, VR designers who wish to use automated testing must recreate the dynamics and graphics of the VR application within the simulation engine used for the biomechanical model. This, however, is impractical for three reasons. First, it introduces implementation overhead, in that two VR environments need to be maintained and updated in parallel. Second, building the replica is often tedious and cumbersome, and requires detailed knowledge of the subtleties of the considered application. Third, the resulting simulation is unlikely to create ecologically valid predictions of interactive user behaviour, as the replica will typically correspond to a low-fidelity version of the original VR application. Such an approach thus considerably complicates alignment of the user input sensed by the application, the (visual) output generated by the application, and the underlying application dynamics themselves. We call this the VR Simulation Alignment Problem and describe it in detail in Section 3.1. Simulations that do not adequately address this problem will provide inferior predictions of how real users interact with VR interfaces or techniques, limiting the validity of in silico testing and compromising the comparative analysis of design choices.
@@
To overcome these limitations, we propose SIM2VR, a system that integrates biomechanical user simulations directly into a given Unity application. Alignment is improved by simulating hardware within the physics engine used for biomechanical simulation, tracking virtual VR input and output using OpenXR, accessing all relevant application data directly in Unity, and synchronising the user simulation and the VR application. From a research perspective, this enables a more accurate and reliable development and validation of user simulation models by ensuring that real and simulated users interact with exactly the same virtual environment. From an interactive software development perspective, our work provides a new solution to the long-standing problem of automated testing of VR interaction.
@@
  • a novel definition and comprehensive discussion of the VR Simulation Alignment Problem encountered when simulating movement-based interaction in VR (Section 3);
  • SIM2VR, a system for integrating biomechanical simulations into Unity applications, which partially addresses the VR Simulation Alignment Problem (Section 4);
  • a demonstration of SIM2VR’s capability to predict differences in performance and ergonomics, and to reveal potential user strategies, verified by an accompanying user study (Section 5).
@@ 2. Background and Context
Below, we situate our work in relation to previous efforts in both automated testing in VR development, and computational user modeling and simulation.
@@ 2.1. VR Development and Automated Testing
Because of the effort and time needed to don and doff equipment during prototype development, iterative testing and development in VR has a higher cost than traditional interface design for desktop or mobile software, making it typically slow and cumbersome (Liu et al., 2023). This underlines the need to automate testing as much as possible. Automated testing allows prototypes to be automatically evaluated and compared without the need for expensive user studies (Bierbaum et al., 2003; Karakaya et al., 2022), by predicting performance estimates such as completion times and error rates.
While the need for automated testing has been clear for decades, a recent study of over 300 VR projects found that 79% of the projects did not use any automatic tests, and that test automation followed common software testing practices without considering the user’s body movements, focusing on aspects such as evaluating the correctness of a response after an event was triggered (Rzig et al., 2023). So far, even automated testing approaches designed specifically for VR do not emulate or simulate the user’s moving body. Instead, they focus on moving and rotating the viewpoint to inspect VR scenes (Wang, 2022), or operate on higher-level action events such as “click UI button” or “grab object” which abstract away the details of the user’s movements (Bierbaum et al., 2003; Harms, 2019; Karakaya et al., 2022). In the domain of Natural User Interfaces (NUIs), a generative motion model was proposed for automatic testing of gestural interaction in (Hunt et al., 2014), but unlike our work, the model was not embedded in a learning loop; instead, it generated random movement sequences without the ability to adapt to the tested interface.
@@ 2.2.1. Simulation in HCI
As discussed in (Murray-Smith et al., 2022), user-centred design could move to a more rigorous, safer and predictable process via a simulation-based approach to interaction design. The same paper highlights general potential benefits of simulation in HCI for efficiency of design, safety, supporting diversity and transparency. Simulations can be used to predict task performance in interaction, such as time taken to finish the task and how often tasks can be successfully accomplished. For example, very early work on keystroke-level models in (Card et al., 1983), and more recent work (Kristensson and Müllners, 2021) used modelling to replace extensive experimentation, optimising text entry system parameters.
The general tools for simulation are improving rapidly, opening up new opportunities. As pointed out in (Lavin et al., 2021), we have seen recent developments in probabilistic, differentiable programming, high-performance computing and causal modelling. This, combined with the rapidly improving ability of machine learning to emulate complex aspects of human perception and behaviour (e.g. deep neural nets can simulate human-level visual perception (He et al., 2016), speech recognition (Graves et al., 2013) and control (Mnih et al., 2015)), means that these new tools for simulation have an increased potential to be efficiently and usefully applied to domains such as HCI.
@@ 2.2.2. Biomechanical User Simulation
Biomechanical models and computer-based simulations are well-established (Ayoub et al., 1974). Increases in computational power have allowed them to evolve from simple models, limited to computation of mechanical loads in static postures (Winter, 1984), to more physiologically-accurate musculoskeletal models (Delp et al., 2007; Damsgaard et al., 2006). Biomechanical models were typically used for inverse simulation, e.g. using the OpenSim ecosystem (Seth et al., 2011). Such inverse simulation methods, namely inverse kinematics, inverse dynamics and static optimisation or computed muscle control, allow the estimation of mechanical loads within the human musculoskeletal system, and neural controls of the muscles given motion tracking data of a given user’s movement as input  (Delp et al., 2007). In HCI, inverse dynamics have been used to identify fundamental ergonomics issues, such as “gorilla arm” (Jang et al., 2017). Forward simulation methods, in contrast, are less frequently used in standalone situations, as they require muscle controls as inputs, which are extremely complex to measure experimentally. However, these have become more useful when used in conjunction with computational controllers (Lee and Umberger, 2016; Dembia et al., 2021; Schumacher et al., 2023) and efficient physics engines (Klar et al., 2023), which we will exploit in this paper.
@@ 2.2.3. Optimal Feedback Control and Reinforcement Learning Models
Control-theoretic models of human motion for mouse movements were developed in (Müller et al., 2017) and supplemented by the assumption of Optimal Feedback Control (Todorov and Jordan, 2002) in (Fischer et al., 2022). In such optimization-based forward simulations, users are assumed to continuously interact with their environment by selecting controls such that a task-specific utility (e.g., a game score) is maximised, given the perceptual information available at decision time. By including intermittent control behaviour typical of humans, simulations could also capture more realistic models of human motor variability (Martín et al., 2021).
Reinforcement learning has recently emerged as a suitable framework for modelling human behaviour in a flexible way: one only needs to define the states, actions and rewards, and then use RL methods to estimate the optimal policy (Chen et al., 2015, 2017). In some cases, the reward function and other parameters can be inferred from human data (Banovic et al., 2016; Kangasrääsiö et al., 2017). If the reward function and state-action space, including their key limitations, are similar to those of a human, increasingly human-like behaviour has been shown to emerge through learning (Oulasvirta et al., 2022).
Recent applications in HCI include models of typing, menu selection, multitasking and visual decision-making (Oulasvirta et al., 2022). Using the assumptions of signal-dependent control noise and movement time minimization, Fischer et al. (Fischer et al., 2021) have shown that an RL agent can generate human-like mid-air trajectories that are in accordance with well-established phenomena such as Fitts’ Law (Fitts, 1954) and the Two-Thirds Power Law (Lacquaniti et al., 1983). RL-based simulation may also provide valuable information for predicting usability- and ergonomics-related criteria and to aid in interface design. For instance, Cheema et al. trained a torque-actuated biomechanical arm model in a mid-air pointing task, and used the model to predict fatigue of real human subjects performing the task (Cheema et al., 2020). In this paper, we build on these developments and use RL to model interaction in VR.
@@ 2.3. Creating Simulated Users
By combining perception models, musculoskeletal models, and physically simulated input devices, we can train RL agents, or “simulated users”, to model and simulate intricate interaction tasks, such as those requiring visuomotor control. An early example of such control was presented in (Nakada et al., 2018), where Nakada et al. introduced a virtual human model and used deep learning to learn reaching and tracking tasks.
@@
The development of computationally efficient simulation software, such as MuJoCo (Todorov et al., 2012), has permitted the use of deep RL for flexible problem formalisations of interaction tasks. This allows a researcher to guide an agent’s learning through reward functions, as was demonstrated in (Ikkala et al., 2022) by Ikkala et al. With their User-in-the-Box (UitB) approach, Ikkala et al. trained a muscle-actuated model of the upper torso and right arm to complete a series of interaction tasks. UitB allows combining biomechanical models with perception models and movement-based interaction tasks through a modular software structure, which facilitates integrating custom model components. In our work, we exploit this modularity and extend UitB to VR environments.
@@
MyoSuite (Caggiano et al., 2022a), another user simulator implemented in MuJoCo, focuses on dexterous hand, lower- and full-body movements. MyoSuite also provides a conversion tool for converting OpenSim models into MuJoCo format (MyoConverter [^3] https://github.com/MyoHub/myoconverter (Wang et al., 2022; Ikkala and Hämäläinen, 2020)), which can also be readily added into the UitB framework.
@@
Simulated users could promote ability-based design in VR, as diverse user populations, i.e., populations with different physical or cognitive properties, could be simulated without the risk of physical harm (Murray-Smith et al., 2022; Quek et al., 2011; Quek, 2013). However, existing approaches that create and evaluate simulated users in the context of HCI are limited to “mock-up” interaction environments that require a re-implementation of the application dynamics real users face during interaction (Fischer et al., 2021; Ikkala et al., 2022; Hetzel et al., 2021). These approaches are thus fundamentally unsuitable for automated biomechanical testing of a given VR prototype. To the best of our knowledge, our work is the first to create simulated users in an ecologically valid VR context by training biomechanical user models directly in the same VR environment that human users interact with.
3. The VR Simulation Alignment Problem
Table 1. Simulated user interaction in VR needs to be aligned with real user interaction in six dimensions. For each dimension, the table provides a description and examples of corresponding variables to be aligned.
Alignment Dimension Description Examples of Alignment Targets User Perceptual Environmental Perception Visual & Proprioceptive Observations Cognitive User Behaviour Muscle Control Signals Physical Biomechanics & Hardware Dynamics Arm Posture, VR Controller Position Appl. Input Channel Input Data Transfer Sensed Position of VR Controller Application VR Application Dynamics Virtual Object Size & Position, Game Score Output Channel Output Data Transfer Rendered Image, Controller Vibrations, Audio Signal
[…]
The table contains the columns ”Alignment Dimension”, ”Description” and ”Examples of Alignment Targets” for the following six rows (the first three corresponding to user-related dimensions, the latter three to application-related dimensions): 1. Perceptual - Environmental Perception - Visual & Proprioceptive Observations. 2. Cognitive - User Behaviour - Muscle Control Signals. 3. Physical - Biomechanics Hardware Dynamics - Arm Posture, VR Controller Position. 4. Input Channel - Input Data Transfer - Sensed Position of VR Controller. 5. Application - VR Application Dynamics - Virtual Object Size Position, Game Score. 6. Output Channel - Output Data Transfer - Rendered Image, Controller Vibrations, Audio Signal.

### v2

@@ Abstract.
Automated biomechanical testing has great potential for the development of VR applications, as initial insights into user behaviour can be gained in silico early in the design process. In particular, it allows prediction of user movements and ergonomic variables, such as fatigue, prior to conducting user studies. However, there is a fundamental disconnect between simulators hosting state-of-the-art biomechanical user models and simulators used to develop and run VR applications. Existing user simulators often struggle to capture the intricacies of real-world VR applications, reducing ecological validity of user predictions. In this paper, we introduce sim2vr, a system that aligns user simulation with a given VR application by establishing a continuous closed loop between the two processes. This, for the first time, enables training simulated users directly in the same VR application that real users interact with. We demonstrate that sim2vr can predict differences in user performance, ergonomics and strategies in a fast-paced, dynamic arcade game. In order to expand the scope of automated biomechanical testing beyond simple visuomotor tasks, advances in cognitive models and reward function design will be needed.
biomechanical simulation, interaction design, VR simulation alignment, automated testing, virtual reality, VR development, VR application, deep reinforcement learning
ccs: Human-centered computing Human computer interaction (HCI)ccs: Human-centered computing Systems and tools for interaction designccs: Human-centered computing Virtual realityccs: Human-centered computing User models
Figure 1. Simulating interactive user behaviour in VR typically requires re-implementing the game dynamics, logic and graphics of the VR application in another simulation engine, which is time-consuming, redundant and error-prone (left).
sim2vr (middle) enables VR designers, for the first time, to run state-of-the-art biomechanical user simulations directly in the same VR application real users are interacting with (right). By aligning simulated and real interaction with respect to VR input and output, as well as application dynamics, sim2vr allows for more ecologically valid predictions of user performance, effort and strategies, supporting developers in the early stages of the design process.
@@ 1. Introduction
Recent advances in deep reinforcement learning (deep RL) and biomechanical models have enabled the creation of simulated users for modelling how humans may behave in interaction tasks (Nakada18; Fischer21; Ikkala22; caggiano2022myosuite). Using advanced biomechanical models that capture the intricacies of human body movement beyond linked-segment models, such simulated users allow prediction of time-continuous movements and effort-related quantities such as muscle energy expenditure (Umberger03) and fatigue (Xia08; Rashedi15). In HCI, simulated users have been recently demonstrated for various tasks, including pointing (Klar22; Cheema20), typing (Hetzel21; Jokinen21), tracking (Fischer21), steering and choice reaction (Ikkala22). This modelling approach is promising, as it enables designers and developers to pose “what-if?” questions to assess how changes in design choices affect users’ performance and ergonomics.
@@
This bears particular promise for HCI domains where human data is scarce or costly to collect, such as in virtual reality (VR) interactions. In VR development, testing prototypes typically involves wearing a headset, standing up and moving around in space, which incurs significant overhead and renders iterative testing slow and cumbersome (liu2023challenges). Biomechanical user simulations show significant potential to remedy this as a means of automated testing. In particular, they allow VR designers to evaluate and compare prototypes without the need for expensive user studies, which can be lengthy and stressful, especially for vulnerable participants. By using an appropriate range of different user models, a simulation-based approach can increase diversity during the testing process, thereby promoting ability-based design (murraysmith2022simulation).
@@
To realise the potential of simulated users, we need to address the challenge of integrating model-based simulations into VR development workflows. State-of-the-art biomechanical models are not readily available in commonly used VR development environments like Unity, [^1] https://unity.com/ or they cannot be combined with deep RL approaches to generate forward simulations in those environments. For instance, while there exists a MuJoCo Unity plugin [^2] https://github.com/google-deepmind/mujoco/tree/main/unity that provides access to biomechanical models implemented in MuJoCo (Todorov12), an efficient physics engine suitable for biomechanical simulation, the models can only be controlled manually using the Unity Editor.
@@
As a result, VR designers who wish to use automated testing must recreate the dynamics and graphics of the VR application within the simulation engine used for the biomechanical model. This, however, is impractical for three reasons. First, it introduces implementation overhead, in that two VR environments need to be maintained and updated in parallel. Second, building the replica is often tedious and cumbersome, and requires detailed knowledge of the subtleties of the considered application. Third, the resulting simulation is unlikely to create ecologically valid predictions of interactive user behaviour, as the replica will typically correspond to a low-fidelity version of the original VR application. Such an approach thus considerably complicates the reduction of differences between the simulated interaction loop and the real interaction loop, which can be described by means of several alignment problems, as detailed in Section 3. Simulations that do not adequately address these problems will provide inferior predictions of how real users interact with VR interfaces or techniques, limiting the validity of in silico testing and compromising the comparative analysis of design choices.
@@
To overcome these limitations, we propose sim2vr, a system that integrates biomechanical user simulations directly into a given Unity application. Alignment is improved by simulating hardware within the physics engine used for biomechanical simulation, tracking virtual VR input and output using OpenXR, accessing all relevant application data directly in Unity, and synchronising the user simulation and the VR application. From a research perspective, this enables a more accurate and reliable development and validation of user simulation models by ensuring that real and simulated users interact with exactly the same virtual environment (see Figure 1). From an interactive software development perspective, our work provides a new solution to the long-standing problem of automated testing of VR interaction.
@@
  • sim2vr, a system that for the first time integrates biomechanical simulations directly into VR application development (Section 4);
  • an empirical assessment of sim2vr’s predictive capabilities for automated testing (Section 5);
  • a critical discussion on the scope as well as current and inherent limitations of sim2vr (Section LABEL:sec:discussion-limitations).
Open-source code of sim2vr is available at https://github.com/fl0fischer/sim2vr.
@@ 2. Background and Context
Below, we situate our work in relation to previous efforts in both automated testing in VR development, and computational user modelling and simulation.
@@ 2.1. VR Development and Automated Testing
Because of the effort and time needed to don and doff equipment during prototype development, iterative testing and development in VR has a higher cost than traditional interface design for desktop or mobile software, making it typically slow and cumbersome (liu2023challenges). This underlines the need to automate testing as much as possible. Automated testing allows prototypes to be automatically evaluated and compared without the need for expensive user studies (bierbaum2003automated; karakaya2022automated), by predicting performance estimates such as completion times and error rates.
While the need for automated testing has been clear for decades, a recent study of over 300 VR projects found that 79% of the projects did not use any automatic tests, and that test automation followed common software testing practices without considering the user’s body movements, focusing on aspects such as evaluating the correctness of a response after an event was triggered (rzig2023virtual). So far, even automated testing approaches designed specifically for VR do not emulate or simulate the user’s moving body. Instead, they focus on moving and rotating the viewpoint to inspect VR scenes (wang2022vrtest), or operate on higher-level action events such as “click UI button” or “grab object” which abstract away the details of the user’s movements (bierbaum2003automated; harms2019automated; karakaya2022automated). In the domain of Natural User Interfaces (NUIs), a generative motion model was proposed for automatic testing of gestural interaction in (hunt2014automatic), but unlike our work, the model was not embedded in a learning loop; instead, it generated random movement sequences without the ability to adapt to the tested interface.
@@ 2.2.1. Simulation in HCI
As discussed in (murraysmith2022simulation), user-centred design could move to a more rigorous, safer and predictable process via a simulation-based approach to interaction design. The same paper highlights general potential benefits of simulation in HCI for efficiency of design, safety, supporting diversity and transparency. Simulations can be used to predict task performance in interaction, such as time taken to finish the task and how often tasks can be successfully accomplished. For example, very early work on keystroke-level models in (card1983psychology), and more recent work (kristensson2021design) used modelling to replace extensive experimentation, optimising text entry system parameters.
The general tools for simulation are improving rapidly, opening up new opportunities. As pointed out in (lavin2021simulation), we have seen recent developments in probabilistic, differentiable programming, high-performance computing and causal modelling. This, combined with the rapidly improving ability of machine learning to emulate complex aspects of human perception and behaviour (e.g. deep neural nets can simulate human-level visual perception (he2016deep), speech recognition (graves2013speech) and control (mnih2015human)), means that these new tools for simulation have an increased potential to be efficiently and usefully applied to domains such as HCI.
@@ 2.2.2. Biomechanical User Simulation
Biomechanical models and computer-based simulations are well-established (Ayoub1974). Increases in computational power have allowed them to evolve from simple models, limited to computation of mechanical loads in static postures (Winter1984), to more physiologically-accurate musculoskeletal models (Delp07; Damsgaard06). Biomechanical models were typically used for inverse simulation, e.g. using the OpenSim ecosystem (SETH2011212). Such inverse simulation methods, namely inverse kinematics, inverse dynamics and static optimisation or computed muscle control, allow the estimation of mechanical loads within the human musculoskeletal system, and neural controls of the muscles given motion tracking data of a given user’s movement as input  (Delp07). In HCI, inverse dynamics have been used to identify fundamental ergonomics issues, such as “gorilla arm” (jang2017modeling). Forward simulation methods, in contrast, are less frequently used in standalone situations, as they require muscle controls as inputs, which are extremely complex to measure experimentally. However, these have become more useful when used in conjunction with computational controllers (lee2016generating; Dembia2021; schumacher2023deprl) and efficient physics engines (Klar22), which we will exploit in this paper.
@@ 2.2.3. Optimal Feedback Control and Reinforcement Learning Models
Control-theoretic models of human motion for mouse movements were developed in (muller2017control) and supplemented by the assumption of Optimal Feedback Control (Todorov02) in (Fischer22). In such optimisation-based forward simulations, users are assumed to continuously interact with their environment by selecting controls such that a task-specific utility (e.g., a game score) is maximised, given the perceptual information available at decision time. By including intermittent control behaviour typical of humans, simulations could also capture more realistic models of human motor variability (martin2021intermittent).
Reinforcement learning has recently emerged as a suitable framework for modelling human behaviour in a flexible way: one only needs to define the states, actions and rewards, and then use RL methods to estimate the optimal policy (chen2015emergence; chen2017cognitive). In some cases, the reward function and other parameters can be inferred from human data (Banovic16; Kangasraasio17). If the reward function and state-action space, including their key limitations, are similar to those of a human, increasingly human-like behaviour has been shown to emerge through learning (oulasvirta2022computational).
Recent applications in HCI include models of typing, menu selection, multitasking and visual decision-making (oulasvirta2022computational). Using the assumptions of signal-dependent control noise and movement time minimisation, Fischer et al. (Fischer21) have shown that an RL agent can generate human-like mid-air trajectories that are in accordance with well-established phenomena such as Fitts’ Law (Fitts54) and the Two-Thirds Power Law (Lacquaniti83). RL-based simulation may also provide valuable information for predicting usability- and ergonomics-related criteria and to aid in interface design. For instance, Cheema et al. trained a torque-actuated biomechanical arm model in a mid-air pointing task, and used the model to predict fatigue of real human subjects performing the task (Cheema20). In this paper, we build on these developments and use RL to model interaction in VR.
@@ 2.3. Creating Simulated Users
By combining perception models, musculoskeletal models, and physically simulated input devices, we can train RL agents, or simulated users, to model and simulate intricate interaction tasks, such as those requiring visuomotor control. An early example of such control was presented in (Nakada18), where Nakada et al. introduced a virtual human model and used deep learning to learn reaching and tracking tasks.
@@
The development of computationally efficient simulation software, such as MuJoCo (Todorov12), has permitted the use of deep RL for flexible problem formalisations of interaction tasks. This allows a researcher to guide an agent’s learning through reward functions, as was demonstrated in (Ikkala22) by Ikkala et al. with their User-in-the-Box (UitB) approach, where they trained a muscle-actuated model of the upper torso and right arm to complete a series of interaction tasks. UitB allows combining biomechanical models with perception models and movement-based interaction tasks through a modular software structure, which facilitates integrating custom model components. In our work, we exploit this modularity and extend UitB to VR environments.
@@
MyoSuite (caggiano2022myosuite), another user simulator implemented in MuJoCo, focuses on dexterous hand, lower- and full-body movements. MyoSuite also provides a conversion tool for converting OpenSim models into MuJoCo format (MyoConverter [^3] https://github.com/MyoHub/myoconverter (wang2022myosim; Ikkala20)), which can also be readily added into the UitB framework.
@@
Simulated users could promote ability-based design in VR, as diverse user populations, i.e., populations with different physical or cognitive properties, could be simulated without the risk of physical harm (murraysmith2022simulation; QueBolWil2011; quek2013phd). However, existing approaches that create and evaluate simulated users in the context of HCI are limited to “mock-up” interaction environments that require a re-implementation of the application dynamics real users face during interaction (Fischer21; Ikkala22; Hetzel21). These approaches are thus fundamentally unsuitable for automated biomechanical testing of a given VR prototype. To the best of our knowledge, our work is the first to create simulated users in an ecologically valid VR context by training biomechanical user models directly in the same VR environment that human users interact with.
3. VR Simulation Alignment